OpenStudy (anonymous):

Which equation has a solution of n = -1 1/2n - 6 =-7 -6n + 8 =-2 2(n-8) = -18 2n + 12 - 3n = 15

OpenStudy (anonymous):

plug in -1 and see whichis correct

OpenStudy (anonymous):

I tried pluggin in but its just so hard to understand

OpenStudy (precal):

let's see -6n+8=2 -6(-1)+8=2 6+8=2 14=2 False, does not work here. Do the rest and post it, we can check your work for you